# BudgetThinker: Empowering Budget-aware LLM Reasoning with Control Tokens

### v1

@@ 1 Introduction
In this paper, we address the challenge of precisely controlling the length of a model’s thought process. We argue that merely stating a budget constraint in the initial prompt is insufficient, and the model must be continuously reminded of its remaining token budget as it generates its response. To achieve this, we develop a novel training and inference framework that enables precise control over the length of the CoT. Our method periodically inserts special Control Tokens that act as explicit signals of the remaining budget, as illustrated in Figure 1a. As shown in Figure 1b, training with these control tokens leads to a faster and more stable reduction in the gap between the generated length and the target budget, demonstrating their effectiveness in teaching budget adherence.
@@ 3.2 Main Performance
Figure 3: Budget Adherence and Utilization Analysis on MATH-500. We compare BudgetThinker-1.5B, ThinkPrune-1.5B, and the original DeepSeek-Distilled-Qwen-2.5-1.5B on MATH-500. (a): The percentage of generated responses that terminate naturally within the specified budget limit ($B$). (b): For responses that finish within the budget, this shows the average generated length as a percentage of the total allocated budget ($|y|/B$). This shows BudgetThinker has a superior ability to follow budget constraints while also making more effective use of the allocated tokens compared to baselines.
@@ 4.2 Efficient Chain of Thoughts
Several methods have been proposed to alleviate the token efficiency problem of LLMs’ reasoning. Prompt methods (Xu et al., 2025b; Han et al., 2024; Muennighoff et al., 2025) make LLMs generate less reasoning tokens directly by adding explicit length constraints into the prompts. Solution routing works (Yu et al., 2025b; Wang et al., 2025a) allow for mid-generation control to prune unpromising traces. Computation routing methods (Damani et al., 2024; Fu et al., 2025; Wang et al., 2025b) choose to allocate just the necessary computation budget based on predicted complexity of queries, either by routing to larger models or conducting more samplings to vote for the final answer. There are also learning methods that either construct well-designed datasets to apply model fine-tuning (Chen et al., 2025b; Han et al., 2024; Xia et al., 2025; Zeng et al., 2025) or use RL methods (Yu et al., 2025a; Yang et al., 2025; Aggarwal & Welleck, 2025; Hou et al., 2025; Arora & Zanette, 2025) to encourage model to generate concise yet accurate answers. More recently, models like Qwen3 (Yang et al., 2025), Kimi k1.5 (Team et al., 2025) and GPT-5 (OpenAI, 2025) integrate hybrid thinking modes of long and short CoTs, seeking trade-off between output length and model performance. Instead, our work achieves precise token budget control. LLMs are trained to understand and follow the constraint in prompts and adaptively adjust their reasoning process to complete the task within the user-requested budgets, showcasing the flexibility and efficiency of our method.
@@ 5 Conclusion
In this work, we introduce BudgetThinker, a novel framework that manages the trade-off between reasoning quality and computational cost in LLMs. By using special control tokens and a two-stage training pipeline, BudgetThinker enables precise, budget-aware control over the model’s reasoning length. Our experiments validate that this method achieves superior budget adherence while maintaining high accuracy on challenging benchmarks. BudgetThinker represents a significant step towards developing more efficient LLMs suitable for real-time, resource-constrained applications.

### v2

@@ 1 Introduction
In this paper, we address the challenge of precisely controlling the length of a model’s thought process. We argue that merely stating a budget constraint in the initial prompt is insufficient, and the model needs to be continuously reminded of its remaining token budget as it generates its response. To achieve this, we develop a novel training and inference framework that enables precise control over the length of the CoT. Our method periodically inserts special Control Tokens that act as explicit signals of the remaining budget, as illustrated in Figure 1a. As shown in Figure 1b, training with these control tokens leads to a faster and more stable reduction in the gap between the generated length and the target budget, demonstrating their effectiveness in teaching budget adherence.
@@ 3.2 Main Performance
Figure 3: Budget following and utilization analysis on MATH-500. We compare BudgetThinker-1.5B, ThinkPrune-1.5B, and the original DeepSeek-Distilled-Qwen-2.5-1.5B on MATH-500. (a): The percentage of generated responses that terminate naturally within the specified budget limit ($B$). (b): For responses that finish within the budget, this shows the average generated length as a percentage of the total allocated budget ($|y|/B$). This shows BudgetThinker has a superior ability to follow budget constraints while also making more effective use of the allocated tokens compared to baselines.
@@ 4.2 Efficient Chain of Thoughts
Several methods have been proposed to alleviate the token efficiency problem of LLMs’ reasoning. Prompt methods (Xu et al., 2025b; Han et al., 2024; Muennighoff et al., 2025) make LLMs generate less reasoning tokens directly by adding explicit length constraints into the prompts. Solution routing works (Yu et al., 2025b; Wang et al., 2025a) allow for mid-generation control to prune unpromising traces. Computation routing methods (Damani et al., 2024; Fu et al., 2025; Wang et al., 2025b; Li et al., 2025) choose to allocate just the necessary computation budget based on predicted complexity of queries, either by routing to larger models or conducting more samplings to vote for the final answer. There are also learning methods that either construct well-designed datasets to apply model fine-tuning (Chen et al., 2025b; Han et al., 2024; Xia et al., 2025; Zeng et al., 2025) or use RL methods (Yu et al., 2025a; Yang et al., 2025; Aggarwal & Welleck, 2025; Hou et al., 2025; Arora & Zanette, 2025) to encourage model to generate concise yet accurate answers. More recently, models like Qwen3 (Yang et al., 2025), Kimi k1.5 (Team et al., 2025) and GPT-5 (OpenAI, 2025) integrate hybrid thinking modes of long and short CoTs, seeking trade-off between output length and model performance. Instead, our work achieves precise token budget control. LLMs are trained to understand and follow the constraint in prompts and adaptively adjust their reasoning process to complete the task within the user-requested budgets, showcasing the flexibility and efficiency of our method.
@@ 5 Conclusion
In this work, we introduce BudgetThinker, a novel framework that manages the trade-off between reasoning quality and computational cost in LLMs. By using special control tokens and a two-stage training pipeline, BudgetThinker enables precise, budget-aware control over the model’s reasoning length. Our experiments validate that this method achieves superior budget adherence while maintaining high accuracy on challenging benchmarks. Thus, BudgetThinker represents a significant step towards developing more efficient LLMs suitable for real-time, resource-constrained applications.